USA Today reported that Parkfield, California, is dubbed the world's earthquake capital because it sits on top of the notorious San Andreas fault. Since 1857, Parkfield has had a major earthquake on the average of 2.1 times every 22 years.

(b) Compute the probability of at least one major earthquake in the next 22 years. Round λ to the nearest hundredth, and use a calculator. (Use 4 decimal places.)

(c) Compute the probability that there will be no major earthquake in the next 22 years. Round λ to the nearest hundredth, and use a calculator. (Use 4 decimal places.)

(d) Compute the probability of at least one major earthquake in the next 46 years. Round λ to the nearest hundredth, and use a calculator. (Use 4 decimal places.)

(e) Compute the probability of no major earthquakes in the next 46 years. Round λ to the nearest hundredth, and use a calculator. (Use 4 decimal places.)
Answer #1

Given that d. 2.1 for 2 years using poisson distai bustion P(x) (6) P(x ) = 1- P(x=0) -1- 2 2.(2.0) = 1-[o. 12246] = (0.8776)